40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 18 17:16:30 2010, Jenne, Update, SUS, Old PRM, SRM stored, new PRM drag wiped 
    Reply  Mon Oct 18 18:00:36 2010, Koji, Update, SUS, Old PRM, SRM stored, new PRM drag wiped 
Message ID: 3737     Entry time: Mon Oct 18 18:00:36 2010     In reply to: 3736
Author: Koji 
Type: Update 
Category: SUS 
Subject: Old PRM, SRM stored, new PRM drag wiped 

- Steve is working on the storage shelf for those optics.

- PRMU002 was chosen as it has the best RoC among the three.

Quote:

[Jenne, Suresh]

We've put the old PRM and SRM (which were living in a foil house on the cleanroom optical table) into Steve's nifty storage containers.  Also, we removed the SRM which was suspended, and stored it in a nifty container.  All 3 of these optics are currently sitting on one of the cleanroom optical tables.  This is fine for temporary storage, but we will need to find another place for them to live permanently.  The etched names of the 3 optics are facing out, so that you can read them without picking them up.  I forgot to note the serial numbers of the optics we've got stored, but the old optics are labeled XRM ###, whereas the new optics are labeled XRMU ###. 

Koji chose for us PRMU 002, out of the set which we recently received from ATF, to be the new PRM.  Suresh and I drag wiped both sides with Acetone and Iso, and it is currently sitting on one of the rings, in the foil house on the cleanroom optical table.

We are now ready to begin the guiderod gluing process (later tonight or tomorrow).

 

ELOG V3.1.3-